OpenStudy (anonymous):

@some1 Indicate in standard form the equation of the line passing through the given points. L(5, 0), M(0, 5)

OpenStudy (anonymous):

okai so do the same thing. x2-x1/y2-y1 5-0/0-5=-1 slope=-1. pick a pair of coordinates, and substitute -1 in for m in y=mx+b y=mx+b 0=-1(5)+b 5=b. y=-x+5. =)

OpenStudy (anonymous):

but..the answer has to start with x_________

OpenStudy (anonymous):

ok then it's x+y=5. Just move the x to the other side. Sorry it's standard form -.- Heh. Ma bad. xD

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how about this one Indicate in standard form the equation of the line passing through the given points. X(0, 6), Y(5, 6) if it starts with y________

OpenStudy (anonymous):

okai. same thing, shell. x2-x1/y2-y1 0-5/6-6 slope=.......uhhhhhhh. undefined...let's just say 0 for now. xD. y=mx+b 6=0(6)+b 6=b. y=6. no that can't be right....o_O

OpenStudy (anonymous):

lol im telling u...its so hard:/

OpenStudy (anonymous):

it's not hard it's just...confoosing. lemme do dis again.

OpenStudy (anonymous):

:/ k

OpenStudy (anonymous):

ok so the slope formula is the other way around xD whoops. but that doesn't change the answer considering we used 0 anyway.....

OpenStudy (anonymous):

ughh darn!!:c

OpenStudy (anonymous):

hm. For the first time. I don't know an algebra question. fml. I had this all down easily. Something's wrong with the problem!!!!! :O

OpenStudy (anonymous):

:c WHYYYYYYY!!

OpenStudy (anonymous):

Yea it's definitely the problem. i just graphed and everything. lol. did you try to put in y=6? does it work?

OpenStudy (anonymous):

it doesnt!!!:c

OpenStudy (anonymous):

Did you copy the problem right into this website? o_O

OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. X(0, 6), Y(5, 6) Place the symbols on the grid to complete the equation. y___________________

OpenStudy (anonymous):

y=6 :D

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

It can't be jolly. The two points share a common coordinate, 6. When a pair share a common coordinate, the equation must be y= or x= that coordinate. lol.

OpenStudy (anonymous):

yea it's y=6. I said that before a;lsdkfj;alskdjf;lasfd

OpenStudy (anonymous):

i ment the x-y=1 for the first question asked. .. might be wrong

OpenStudy (anonymous):

omg :c this is soo hard!!!!!!!!! :c

OpenStudy (anonymous):

and this one? Indicate the equation of the given line in standard form. The line containing the longer diagonal of a quadrilateral whose vertices are A (2, 2), B(-2, -2), C(1, -1), and D(6, 4).

OpenStudy (anonymous):

what about that question?

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

you on the same lesson?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

what question is it?

OpenStudy (anonymous):

this one: Indicate the equation of the given line in standard form. The line containing the longer diagonal of a quadrilateral whose vertices are A (2, 2), B(-2, -2), C(1, -1), and D(6, 4). xDDD

OpenStudy (anonymous):

lol i ment.. what numer @jolly

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

@some1 you do AOA too? i'm on 16 section 2

OpenStudy (anonymous):

lol i have no idea buddy. But list problems and i'll solve em for ya. :P

OpenStudy (anonymous):

we're not on the same lesson then.. cause 16 is this one for me ->Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of

OpenStudy (anonymous):

woops. i meant #22.

OpenStudy (anonymous):

ohhhhh:)

OpenStudy (anonymous):

hey @some1 for this question: Indicate the equation of the given line in standard form. The line containing the longer diagonal of a quadrilateral whose vertices are A (2, 2), B(-2, -2), C(1, -1), and D(6, 4). is it x=y? or did i do it wrong?

OpenStudy (anonymous):

i know BD is the longer diagonal

OpenStudy (anonymous):

6+2/4+2 8/6 would be slope. check your double negatives. they become positive. like in english, shell! :D

OpenStudy (anonymous):

:D awesome!

OpenStudy (anonymous):

so 4/3 simplified, and y+2 = 4/3(x+2)?

OpenStudy (anonymous):

yah.

OpenStudy (anonymous):

y = 4x +6 is what i got ..

OpenStudy (anonymous):

if u simplify that's what u should get =D

OpenStudy (anonymous):

k, thanks :)